# Zephyrus: Scaling Gateways Beyond the Petabit-Era with DPU-Augmented Hierarchical Co-Offloading

### v1

@@ Abstract
footnotetext: The first three authors contribute equally. Tong Yang (yangtong@pku.edu.cn) and Xiyun Xu (wumin.stephen@bytedance.com) are corresponding authors.
1 Introduction
@@
As shown in Figure 1, ByteDance’s cloud and edge services generate petabit-scale aggregate traffic across tens of thousands of tenants. Orchestrating these workloads critically depends on cloud gateways, which sit at aggregation points and mediate intra-datacenter (east-west) and external (north-south) traffic [1, 2], including IDC and cross-region connectivity. As the linchpin of our network, gateways must simultaneously deliver low latency, sustain massive throughput, and remain cost efficient at scale [3, 4]. These requirements make gateway one of the most performance-critical components in our infrastructure.
@@
Diverse workloads drive the demand for high-performance gateways. On our cross-region LLM platform Doubao [5], user requests are first processed at nearby edge PoPs and then traverse backbone links to regional clusters, generating sustained cross-domain traffic through border gateways. On the short-video platform TikTok [6], video clips are relayed across edge servers while real-time AI features (e.g., filters, face tracking, speech recognition) run at the edge and heavier tasks are handled in the cloud; this hybrid execution creates continuous, high-bandwidth flows between edge and cloud. On Volcengine [7], our enterprise cloud platform, multi-region deployments synchronize storage and services across sites, driving frequent large-scale cross-region transfers, especially in hybrid on-premises–cloud settings. Together, these workloads place gateways on the critical path, requiring them to sustain massive throughput, low-latency forwarding, and scalable multi-tenant policies.
@@
Figure 1: Aggregate traffic from ByteDance cloud and edge services
@@
Our experience with two generations of cloud gateways, together with observations from prior deployments and research, highlights the fundamental limitations of existing approaches. Our first generation of CPU-based software gateways provides flexibility, but fails to scale with exponential traffic growth, suffering from per-core bottlenecks, performance isolation issues, and vulnerability to flow imbalance [8, 9, 10]. The second version of Tofino-based gateways delivers line-rate forwarding with microsecond latency, but faces hard resource ceilings. Although algorithmic LPM (ALPM) [11, 12] and pipeline folding [4, 13] can be applied to expand capacity, production workloads still induced severe localized hotspots. Other works also explore the possibility of FPGA-based gateways, which offer greater flexibility and can implement relatively complex operations that exceed the limits of ASIC resources [3]. However, FPGAs incur substantially higher hardware cost and power consumption compared to programmable ASICs [14, 15]. Moreover, FPGA programming still relies on low-level HDLs (e.g., Verilog, VHDL) and complex vendor toolchains, resulting in much longer development cycles and higher overheads [16, 17, 18]. Taken collectively, these limitations underscore the fundamental tension between performance, programmability, development difficulty, and cost efficiency in gateway design.
@@
However, designing and deploying Zephyrus is a non-trivial task. First, Zephyrus requires precise coordination across heterogeneous components, where task partitioning, functional integration, effective communication between ASICs, DPUs, and software must be carefully orchestrated. Second, DPUs are relatively unexplored in the context of cloud gateways, which compelled us to address their integration complexity largely without existing design blueprints  [3]. Last but not least, system upgrades at petabit scale require careful planning to ensure business continuity, smooth expansion, and robust disaster recovery.
To this end, we design a novel Hierarchical Co-Offloading (HLCO) architecture, which extends the classic fast/slow path model into three tiers: Switching ASICs handle high-throughput, low-complexity processing; DPUs provide a middle tier with greater flexibility; and software offers full programmability for most flexible processing, allowing Zephyrus to harness the strengths of heterogeneous hardware while mitigating their weaknesses. For compatibility, we implemented Zephyrus on P4-programmable DPUs (Pensando [19]) and switching ASICs (Tofino), as both support P4 and thus enable a unified programming model. Four Pensando DPUs are integrated with a folded Tofino pipeline in our deployment, achieving 1.6 Tbps line-rate forwarding. To support efficient coordination, we carefully design a minimal yet expressive metadata format for communication between Tofinos and Pensando DPUs. In the control plane, we develop a version-based flow update mechanism to ensure synchronization of the forwarding rules. Beyond introducing a new gateway design, this paper shares the experiences and lessons distilled from several years of development and production deployment, providing insights that may inform both researchers and practitioners building next-generation cloud gateways.
@@
  • We propose a novel hierarchically co-offloading architecture (HLCO) that enables efficient partitioning of P4 data-plane processing across high-performance switching ASICs and DPUs. HLCO combines the strength of the classic fast/slow path model and flexible 3-stage offloading [20], allowing Zephyrus to harness the strengths of heterogeneous hardware while mitigating their weaknesses. With HLCO, more than 99% of packets are offloaded to hardware, minimizing software overhead while retaining the flexibility.
  • We evaluate Zephyrus against LuoShen [4], a state-of-the-art programmable gateway. Zephyrus delivers 33% higher throughput and is estimated to reduce power consumption by at least 21% and hardware procurement cost by 14% per 2U chassis, under comparable configurations. Compared to Albatross [3], another state-of-the-art FPGA-based cloud gateway, Zephyrus achieves 2× higher throughput while requiring significantly lower estimated total cost of ownership (TCO), demonstrating a superior performance-per-dollar ratio.
2 Background and Motivation
Figure 2: Cloud network gateway architecture.
@@ 2 Background and Motivation
2.1 Anatomy of Cloud Gateways
Roles of cloud gateways. As shown in Figure 2, the Cloud Gateway Architecture across Regions, cloud gateways are packet forwarding devices at the network boundary, serving as the linchpin of the cloud network by orchestrating all data flows [21]. As the central hub, they manage major traffic patterns from east to west (VM-VM) to north to south (VM-cross-region-VM and VM-IDC) [22, 23] with their corresponding datapaths detailed in Table 1. In ByteDance, cloud gateways serve as the high-performance backbone for a wide range of upper-layer applications (e.g., distributed AI services, large-scale content delivery, and hybrid cloud interconnection), providing the low-latency, high-throughput forwarding required for their business-critical traffic.
Table 1: Major cloud services and corresponding traffic routes (CGW for could gateway)
@@ 2.1 Anatomy of Cloud Gateways
Characteristics of gateway workloads. Gateway workloads exhibit several important characteristics. (1) A significant portion of traffic is highly latency-sensitive [24, 25]. As mentioned above, critical applications such as LLM inference, real-time collaboration, and recommendation services impose strict latency limits, where even minor delays can degrade service quality and ultimately compromise service level objectives (SLO). (2) The traffic is characterized by its immense aggregate volume, with aggregate traffic reaching the petabit-per-second (Pbps) scale. This is a direct consequence of simultaneously serving numerous data-intensive applications, from large-scale content delivery to distributed AI services. (3) The traffic stream is inherently a composite of millions of concurrent flows [26, 27]. This high degree of concurrency arises from the gateway’s role as a nexus for a multi-tenant environment, where countless independent application sessions are active simultaneously. (4) Traffic exhibits a highly skewed flow size distribution [28], characterized by the presence of long-lived, high-volume elephant flows alongside a large number of short-lived mice flows. (5) Processing logic is relatively simple and fixed [29], primarily driven by recurring functions such as ACL checks, metering, and next-hop determination, rather than arbitrary per-flow computations. It is also largely stateless in nature, with most flows requiring only basic forwarding or simple policy enforcement.
@@
2.2 Evolution towards Zephyrus
@@
The first generation: Software-only cloud gateways. Our initial cloud gateway adopted a fully software-based architecture, which is flexible and versatile but ultimately limited in performance and scalability [30, 31, 32]. This choice, driven by the need for rapid feature evolution under an MVP strategy, lets us leverage mature user-space packet I/O ecosystems (e.g., DPDK [33], netmap [34], and IX [35]) for fast development. However, it embodied a fundamental contradiction between general-purpose CPUs for diverse computations and the repetitive, structurally simple work of packet forwarding  [36, 37].
As traffic grew, tenant performance isolation became critical. While a default per-attachment cap (e.g., 5 Gbps) effectively contained noisy neighbors, this rigid approach inherently conflicted with east-west traffic demands, resulting in non-negligible packet drops and application degradation. To mitigate this, we explored edge-based congestion-aware limiters (e.g., EyeQ [38]). However, their susceptibility to non-stationary signals and delayed feedback induces oscillatory throttling under bursts, degrading tail latency by a factor of 1.7–4.9 in production [39]. Perhaps the most critical bottleneck, even with horizontal scaling, is the elephant flow problem. Hash-based load balancers, from in-network ECMP to host-level RSS, deterministically map large flows to a single CPU core. This creates severe hotspots that fundamentally limit the entire system’s throughput to that of its single most overloaded core [40]. This per-core bottleneck is a fundamental barrier to scalability, as simply adding more servers fails to address the underlying traffic skew.
@@
The second generation: Tofino-based gateway with x86 CPU assist. To address the challenges inherent to software-based gateways, we designed Tofino-based cloud gateways with the assist of x86 CPUs. Using programmable ASICs, our cloud gateways achieve line-rate performance, delivering 1.6 Tbps of throughput per device while maintaining a consistent ultra-low latency of 2µs. However, even after integrating four independent pipelines into a unified processing unit, the Tofino ASIC suffers from critically scarce on-chip resources. These constraints make supporting large-scale table entries and somewhat complex processing logic extremely challenging. As our tenants continue to grow, the on-chip SRAM required to store massive routing tables (most notably the VM-to-Node-Controller mappings) becomes nearly exhausted, despite applying optimizations including compiler-driven resource packing[41] and key-compression for IPv6 entries[13]. Table 2 details the utilization of SRAM in twelve consecutive pipeline stages within the fourth pipeline where the P4 compiler maps the VM-NC tables, a position rigidly dictated by the data flow dependencies of the previous processing stages.
Table 2: SRAM Utilization Hotspot in Pipeline3. The extremely high average of 97.4% across 12 stages, with five individual stages hitting the 100.0% physical limit, reveals a severe and unavoidable localized resource bottleneck.
@@
In addition to localized on-chip memory shortage, offloading the data plane to Tofinos significantly reduces development flexibility. Some complex operations are difficult to implement efficiently on Tofino pipelines (e.g., NAT, sophisticated Layer-4 load balancing [42, 43]) . Furthermore, computationally intensive functions such as IPsec encryption or operations requiring large elastic buffers, such as IP fragment reassembly, are fundamentally ill-suited for the relatively simple and fixed processing model of Tofinos, forcing these tasks back to software [44].
@@
  • Lower development complexity and faster deployment. Compared to FPGAs, DPUs are significantly easier to develop with. DPU programming typically relies on high-level languages such as C/C++ and P4, supported by mature software ecosystems and toolchains [45, 46]. In contrast, FPGA development requires low-level hardware description languages (e.g.,Verilog / VHDL) and involves long synthesis and verification cycles [47, 48]. These differences make DPUs far more developer-friendly.
  • Easier software access. DPUs integrate on-board ARM cores, enabling native DPDK deployment without external dependencies, as the ARM execution environment supports direct DPDK execution [46, 45]. In contrast, most FPGAs rely on host CPUs via PCIe interconnects, which require custom drivers and introduce complexity for DPDK integration [49, 50]. This architectural difference simplifies DPDK implementation on DPUs by eliminating host-dependent communication layers.
@@ 3.1 Guiding Principles
3.2 High-Level Architecture
Figure 3: Three typical datapaths for different types of traffic
@@ 3.2 High-Level Architecture
Figure 4: Detailed pipeline organization.
Figure 5: Table lookup sequence in Tofino pipelines.
4 Compositional P4 Pipeline Design
@@
4.1 Detailed Pipeline Organization
@@
We also evaluated an alternative design where the DPUs are placed serially after the entire folded Tofino pipeline. Although this arrangement might appear to simplify management, it introduces two critical drawbacks. First and most fundamentally, it misaligns the roles of the hardware. This design forces DPUs to handle line-rate forwarding, a task for which they are not optimized. DPUs excel at offloading complex operational tasks (e.g., encryption, deep packet inspection), not replacing the raw forwarding power of a switching ASIC [51]. Second, the alternative design suffers from a severe bandwidth penalty. This forces a hard partitioning of the front-panel interfaces that half of the ports to be consumed by the internal ASIC-DPU connection and the other half for external connectivity. Consequently, the bandwidth is halved from 1.6 Tbps to 800 Gbps. The key advantage of the original design is that traffic enters and exits the DPU from the same Tofino pipeline, achieving full-duplex communication without bandwidth loss, and it also places the DPU exactly where complex operations are needed. By comparison, LuoShen sacrifices a quarter of its external ports to connect with the CPU and FPGA [4], while the current design avoids any bandwidth loss by allowing data to enter and exit the DPU from the same Tofino pipeline. Therefore, we adopt the three-stage "sandwich" architecture for our design.
Figure 6: DPU packet processing pipeline with fast/slow path interaction
4.2 Packet Processing across Heterogeneous Hardware

### v2

@@ Abstract.
conference: ; ; footnotetext: The first three authors contribute equally. Tong Yang (yangtong@pku.edu.cn) and Xiyun Xu (wumin.stephen@bytedance.com) are corresponding authors.
1. Introduction
@@
As shown in Figure 1, ByteDance’s cloud and edge services generate petabit-scale aggregate traffic across tens of thousands of tenants. Orchestrating these workloads critically depends on cloud gateways, which sit at aggregation points and mediate intra-datacenter (east-west) and external (north-south) traffic (google:jupiter; google:maglev), including IDC and cross-region connectivity. As the linchpin of our network, gateways must simultaneously deliver low latency, sustain massive throughput, and remain cost efficient at scale (lu2025albatross; pan2024luoshen). These requirements make gateway one of the most performance-critical components in our infrastructure.
@@
Diverse workloads drive the demand for high-performance gateways. On our cross-region LLM platform Doubao (doubao), user requests are first processed at nearby edge PoPs and then traverse backbone links to regional clusters, generating sustained cross-domain traffic through border gateways. On the short-video platform TikTok (tiktok), video clips are relayed across edge servers while real-time AI features (e.g., filters, face tracking, speech recognition) run at the edge and heavier tasks are handled in the cloud; this hybrid execution creates continuous, high-bandwidth flows between edge and cloud. On Volcengine (volcengine), our enterprise cloud platform, multi-region deployments synchronize storage and services across sites, driving frequent large-scale cross-region transfers, especially in hybrid on-premises–cloud settings. Together, these workloads place gateways on the critical path, requiring them to sustain massive throughput, low-latency forwarding, and scalable multi-tenant policies.
@@
Figure 1. Aggregate traffic from ByteDance cloud and edge services
@@
Our experience with two generations of cloud gateways, together with observations from prior deployments and research, highlights the fundamental limitations of existing approaches. Our first generation of CPU-based software gateways provides flexibility, but fails to scale with exponential traffic growth, suffering from per-core bottlenecks, performance isolation issues, and vulnerability to flow imbalance (lei2024enhancing; pismenny2021alleviating; kfoury2024smartnic). The second version of Tofino-based gateways delivers line-rate forwarding with microsecond latency, but faces hard resource ceilings. Although algorithmic LPM (ALPM) and pipeline folding (pan2024luoshen; pan2021sailfish) can be applied to expand capacity, production workloads still induced severe localized hotspots. Other works also explore the possibility of FPGA-based gateways, which offer greater flexibility and can implement relatively complex operations that exceed the limits of ASIC resources (lu2025albatross). However, FPGAs incur substantially higher hardware cost and power consumption compared to programmable ASICs (wang2020multitenancy; yang2024p4eip). Moreover, FPGA programming still relies on low-level HDLs (e.g., Verilog, VHDL) and complex vendor toolchains, resulting in much longer development cycles and higher overheads (sozzo2022survey; jamal2018fpga; windh2015fpga). Taken collectively, these limitations underscore the fundamental tension between performance, programmability, development difficulty, and cost efficiency in gateway design.
@@
However, designing and deploying Zephyrus is a non-trivial task. First, Zephyrus requires precise coordination across heterogeneous components, where task partitioning, functional integration, effective communication between ASICs, DPUs, and software must be carefully orchestrated. Second, DPUs are relatively unexplored in the context of cloud gateways, which compelled us to address their integration complexity largely without existing design blueprints  (lu2025albatross). Last but not least, system upgrades at petabit scale require careful planning to ensure business continuity, smooth expansion, and robust disaster recovery.
To this end, we design a novel Hierarchical Co-Offloading (HLCO) architecture, which extends the classic fast/slow path model into three tiers: Switching ASICs handle high-throughput, low-complexity processing; DPUs provide a middle tier with greater flexibility; and software offers full programmability for most flexible processing, allowing Zephyrus to harness the strengths of heterogeneous hardware while mitigating their weaknesses. For compatibility, we implemented Zephyrus on P4-programmable DPUs (Pensando (pensando2022)) and switching ASICs (Tofino), as both support P4 and thus enable a unified programming model. Four Pensando DPUs are integrated with a folded Tofino pipeline in our deployment, achieving 1.6 Tbps line-rate forwarding. To support efficient coordination, we carefully design a minimal yet expressive metadata format for communication between Tofinos and Pensando DPUs. In the control plane, we develop a version-based flow update mechanism to ensure synchronization of the forwarding rules. Beyond introducing a new gateway design, this paper shares the experiences and lessons distilled from several years of development and production deployment, providing insights that may inform both researchers and practitioners building next-generation cloud gateways.
@@
  • We propose a novel hierarchically co-offloading architecture (HLCO) that enables efficient partitioning of P4 data-plane processing across high-performance switching ASICs and DPUs. HLCO combines the strength of the classic fast/slow path model and flexible 3-stage offloading (li2024triton), allowing Zephyrus to harness the strengths of heterogeneous hardware while mitigating their weaknesses. With HLCO, more than 99% of packets are offloaded to hardware, minimizing software overhead while retaining the flexibility.
  • We evaluate Zephyrus against LuoShen (pan2024luoshen), a state-of-the-art programmable gateway. Zephyrus delivers 33% higher throughput and is estimated to reduce power consumption by at least 21% and hardware procurement cost by 14% per 2U chassis, under comparable configurations. Compared to Albatross (lu2025albatross), another state-of-the-art FPGA-based cloud gateway, Zephyrus achieves 2× higher throughput while requiring significantly lower estimated total cost of ownership (TCO), demonstrating a superior performance-per-dollar ratio.
2. Background and Motivation
Figure 2. Cloud network gateway architecture.
@@ 2. Background and Motivation
2.1. Anatomy of Cloud Gateways
Roles of cloud gateways. As shown in Figure 2, the Cloud Gateway Architecture across Regions, cloud gateways are packet forwarding devices at the network boundary, serving as the linchpin of the cloud network by orchestrating all data flows (chiang2018andromeda). As the central hub, they manage major traffic patterns from east to west (VM-VM) to north to south (VM-cross-region-VM and VM-IDC) (greenberg2009vl2; dean2013tail) with their corresponding datapaths detailed in Table 1. In ByteDance, cloud gateways serve as the high-performance backbone for a wide range of upper-layer applications (e.g., distributed AI services, large-scale content delivery, and hybrid cloud interconnection), providing the low-latency, high-throughput forwarding required for their business-critical traffic.
Table 1. Major cloud services and corresponding traffic routes (CGW for could gateway)
@@ 2.1. Anatomy of Cloud Gateways
Characteristics of gateway workloads. Gateway workloads exhibit several important characteristics. (1) A significant portion of traffic is highly latency-sensitive (barker2010empirical; wu2023optimizing). As mentioned above, critical applications such as LLM inference, real-time collaboration, and recommendation services impose strict latency limits, where even minor delays can degrade service quality and ultimately compromise service level objectives (SLO). (2) The traffic is characterized by its immense aggregate volume, with aggregate traffic reaching the petabit-per-second (Pbps) scale. This is a direct consequence of simultaneously serving numerous data-intensive applications, from large-scale content delivery to distributed AI services. (3) The traffic stream is inherently a composite of millions of concurrent flows (chen2014concurrent; bates2014detecting). This high degree of concurrency arises from the gateway’s role as a nexus for a multi-tenant environment, where countless independent application sessions are active simultaneously. (4) Traffic exhibits a highly skewed flow size distribution (pasham2022graph), characterized by the presence of long-lived, high-volume elephant flows alongside a large number of short-lived mice flows. (5) Processing logic is relatively simple and fixed (springCloudGateway), primarily driven by recurring functions such as ACL checks, metering, and next-hop determination, rather than arbitrary per-flow computations. It is also largely stateless in nature, with most flows requiring only basic forwarding or simple policy enforcement.
@@
2.2. Evolution towards Zephyrus
@@
The first generation: Software-only cloud gateways. Our initial cloud gateway adopted a fully software-based architecture, which is flexible and versatile but ultimately limited in performance and scalability (Kohler:Click:SOSP1999; Dobrescu:RouteBricks:SOSP2009; Han:PacketShader:SIGCOMM2010). This choice, driven by the need for rapid feature evolution under an MVP strategy, lets us leverage mature user-space packet I/O ecosystems (e.g., DPDK (dpdk_on_dpu), netmap (Rizzo:netmap:ATC2012), and IX (Belay:IX:OSDI2014)) for fast development. However, it embodied a fundamental contradiction between general-purpose CPUs for diverse computations and the repetitive, structurally simple work of packet forwarding.
As traffic grew, tenant performance isolation became critical. While a default per-attachment cap (e.g., 5 Gbps) effectively contained noisy neighbors, this rigid approach inherently conflicted with east-west traffic demands, resulting in non-negligible packet drops and application degradation. To mitigate this, we explored edge-based congestion-aware limiters (e.g., EyeQ (EyeQ-NSDI13)). However, their susceptibility to non-stationary signals and delayed feedback induces oscillatory throttling under bursts, degrading tail latency by a factor of 1.7–4.9 in production (Protego-NSDI23). Perhaps the most critical bottleneck, even with horizontal scaling, is the elephant flow problem. Hash-based load balancers, from in-network ECMP to host-level RSS, deterministically map large flows to a single CPU core. This creates severe hotspots that fundamentally limit the entire system’s throughput to that of its single most overloaded core (Hedera-NSDI10). This per-core bottleneck is a fundamental barrier to scalability, as simply adding more servers fails to address the underlying traffic skew.
@@
The second generation: Tofino-based gateway with x86 CPU assist. To address the challenges inherent to software-based gateways, we designed Tofino-based cloud gateways with the assist of x86 CPUs. Using programmable ASICs, our cloud gateways achieve line-rate performance, delivering 1.6 Tbps of throughput per device while maintaining a consistent ultra-low latency of 2µs. However, even after integrating four independent pipelines into a unified processing unit, the Tofino ASIC suffers from critically scarce on-chip resources. These constraints make supporting large-scale table entries and somewhat complex processing logic extremely challenging. As our tenants continue to grow, the on-chip SRAM required to store massive routing tables (most notably the VM-to-Node-Controller mappings) becomes nearly exhausted, despite applying optimizations including compiler-driven resource packing and key-compression for IPv6 entries(pan2021sailfish). Table 2 details the utilization of SRAM in twelve consecutive pipeline stages within the fourth pipeline where the P4 compiler maps the VM-NC tables, a position rigidly dictated by the data flow dependencies of the previous processing stages.
Table 2. SRAM Utilization Hotspot in Pipeline3. The extremely high average of 97.4% across 12 stages, with five individual stages hitting the 100.0% physical limit, reveals a severe and unavoidable localized resource bottleneck.
@@
In addition to localized on-chip memory shortage, offloading the data plane to Tofinos significantly reduces development flexibility. Some complex operations are difficult to implement efficiently on Tofino pipelines (e.g., NAT, sophisticated Layer-4 load balancing) . Furthermore, computationally intensive functions such as IPsec encryption or operations requiring large elastic buffers, such as IP fragment reassembly, are fundamentally ill-suited for the relatively simple and fixed processing model of Tofinos, forcing these tasks back to software.
@@
  • Lower development complexity and faster deployment. Compared to FPGAs, DPUs are significantly easier to develop with. DPU programming typically relies on high-level languages such as C/C++ and P4, supported by mature software ecosystems and toolchains (dpdk_on_dpu; nvidia_dpu_whitepaper). In contrast, FPGA development requires low-level hardware description languages (e.g.,Verilog / VHDL) and involves long synthesis and verification cycles (xilinx_hdl). These differences make DPUs far more developer-friendly.
  • Easier software access. DPUs integrate on-board ARM cores, enabling native DPDK deployment without external dependencies, as the ARM execution environment supports direct DPDK execution (nvidia_dpu_whitepaper; dpdk_on_dpu). In contrast, most FPGAs rely on host CPUs via PCIe interconnects, which require custom drivers and introduce complexity for DPDK integration (xilinx_smartnic). This architectural difference simplifies DPDK implementation on DPUs by eliminating host-dependent communication layers.
@@ 3.1. Guiding Principles
3.2. High-Level Architecture
Figure 3. Three typical datapaths for different types of traffic
@@ 3.2. High-Level Architecture
Figure 4. Detailed pipeline organization.
Figure 5. Table lookup sequence in Tofino pipelines.
4. Compositional P4 Pipeline Design
@@
4.1. Detailed Pipeline Organization
@@
We also evaluated an alternative design where the DPUs are placed serially after the entire folded Tofino pipeline. Although this arrangement might appear to simplify management, it introduces two critical drawbacks. First and most fundamentally, it misaligns the roles of the hardware. This design forces DPUs to handle line-rate forwarding, a task for which they are not optimized. DPUs excel at offloading complex operational tasks (e.g., encryption, deep packet inspection), not replacing the raw forwarding power of a switching ASIC (ermerins2021bluefield). Second, the alternative design suffers from a severe bandwidth penalty. This forces a hard partitioning of the front-panel interfaces that half of the ports to be consumed by the internal ASIC-DPU connection and the other half for external connectivity. Consequently, the bandwidth is halved from 1.6 Tbps to 800 Gbps. The key advantage of the original design is that traffic enters and exits the DPU from the same Tofino pipeline, achieving full-duplex communication without bandwidth loss, and it also places the DPU exactly where complex operations are needed. By comparison, LuoShen sacrifices a quarter of its external ports to connect with the CPU and FPGA (pan2024luoshen), while the current design avoids any bandwidth loss by allowing data to enter and exit the DPU from the same Tofino pipeline. Therefore, we adopt the three-stage ”sandwich” architecture for our design.
Figure 6. DPU packet processing pipeline with fast/slow path interaction
4.2. Packet Processing across Heterogeneous Hardware